A non-ideal 1-kg ball is 10-m above a table when it is dropped. It bounces to a height of 7-m.


a) How much energy is transferred to heat & sound during the bounce?


 



b) Explain why this ball cannot bounce to a height of 12-m if it is dropped.





c) What could you do to make the ball bounce to a height of 12-m?


1
Expert's answer
2021-03-04T11:55:24-0500

a) The initial energy of the ball is (potential energy):


"E_1 = mgh_1"

where "m = 1kg" is the mass of the ball, "g = 9.8N\/kg" is the gravitational acceleration, and "h_1 = 10m" is the initial height.

The energy of the ball after it bounces to a height of "h_2 = 7m" is:


"E_2 = mgh_2"

The amount of energy loss is:


"Q = E_1 - E_2 = mg(h_1 - h_2)\\
Q = 1\cdot 9.8\cdot (10-7) = 29.4J"

b) According to the energy conservation law, the ball can not gain extra energy from nothing to bounce to a height of 12-m. It can only bounce not higher then initial 10 m.


c) We can give extra energy to the ball. For example, not releasing it at the beginning, but throwing it downward.